George invested some money at 5% interest. George also invested $165 more than 5 times that amount at 12%. How much is invested at each rate if George receives $899.90 in interest after one year?
Answer by scott8148(5895)

let x="amount @ 5%", so 5x+165="amount @ 12%"

(5%)(x)+(12%)(5x+165)=899.9 __ .05x+.6x+19.8=899.9

.65x=880.1